Joint Cost Allocation—Market Value at Split-off Method

Sugar Sweetheart, Inc., jointly produces raw sugar, granulated sugar, and caster sugar. After the split-off point, raw sugar is immediately sold for $0.20 per pound, while granulated and caster sugar are processed further. The market value of the granulated sugar and caster sugar is estimated to both be $0.25 at the split-off point. One batch of joint production costs $1,640 and yields 3,000 pounds of raw sugar, 3,600 pounds of granulated sugar, and 2,000 pounds of caster sugar at the split-off point.

Allocate the joint costs of production to each product using the market value at split-off method.

Joint Product Allocation
Raw sugar $
Granulated sugar
Caster sugar
Totals $
Answer #1

Joint production costs = $1,640

Fully Processed Monthly
Unit Output

Sales Price
per Unit

Net Realizable
Value

Percent of
Value

Estimated Allocated
Joint Costs

Raw sugar

3,000

$0.20

$600

600/2,000 = 30

%

1,640 X 30% = $492

Granulated sugar

3,600 $0.25 $900 900/2,000 = 45

%

1,640 X 45% = $738

Caster sugar

2,000 $0.25 $500 500/2,000 = 25

%

1,640 x 25% = $410

Total

$2,000

100

%

$1,640

Joint Product Allocation
Raw sugar $492
Granulated sugar $738
Caster sugar $410
Totals $1,640

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.